Proof whether the following operations are inner product operations:

⟨x, y⟩ = 2x1y1 − x1y2 − x2y1 + 2x2y2, x=(x1, x2), y=(y1, y2)



1
Expert's answer
2021-11-21T13:35:18-0500

1. Positive Definite Property:


"\langle x,x\rangle=2x_1x_1-x_1x_2-x_2x_1+2x_2x_2"

"=(x_1^2+x_2^2)+(x_1-x_2)^2\geq0"

The value equals zero if and only if both summands are zero, i.e., when "x_1=x_2=0"

For any "x\in V, \langle x,x\rangle\geq0;" and "\langle x,x\rangle=0" if and only if "x=0."


2. Symmetric Property


"\langle x,y\rangle=2x_1y_1-x_1y_2-x_2y_1+2x_2y_2"

"=2y_1x_1-y_1x_2-y_2x_1+2y_2x_2=\langle y, x\rangle"

3. Linearity


"\langle ax+by,z\rangle"

"=2(ax_1+by_1)z_1-(ax_1+by_1)z_2-(ax_2+by_2)z_1"

"+2(ax_2+by_2)z_2"

"=a(2x_1z_1-x_1z_2-x_2z_1+x_2z_2)"

"+b(2y_1z_1-y_1z_2-y_2z_1+y_2z_2)"

"=a\langle x,z\rangle+b\langle y,z\rangle"

Then "\langle x,y\rangle" is an inner product on "V."